OpenStudy (anonymous):

a=9/2(x+23) solve for x

OpenStudy (anonymous):

a= 9/2x + 207/2 a - 207/2 = 9/2x (a-207/2) / (9/2) = x

OpenStudy (anonymous):

Thx